# FC–PINO: High-Precision Physics-Informed Neural Operators via Fourier Continuation

### v2

@@ 4.2.2 Results
We compare the performance of FC–PINO to the baseline models for solving a single instance of the 2D Burgers’ equation, where derivatives in the physics loss are computed using spectral differentiation. For each architecture, we performed a hyperparameter grid search, varying the number of Fourier modes, learning rate, scheduler patience, and the $(c,d)$ continuation parameters. The optimal hyperparameters for each architecture are provided in Section˜C.2, while Table˜3 reports the errors achieved by the best model for each architecture. The results show that the FC–PINO architecture performs very well, surpassing Pad–PINO by two orders of magnitude, while PINO Out–Pad and Standard PINO again fail to satisfy the governing equations.
@@ 4.3.2 Results
We compare the performance of FC–PINO to the baseline models for solving a single instance of the the Navier–Stokes equations, where derivatives in the physics loss are computed using spectral differentiation. For each architecture, we performed a hyperparameter grid search, varying the number of Fourier modes, learning rate, scheduler patience, and the $(c,d)$ continuation parameters. The optimal hyperparameters for each architecture are provided in Section˜C.3, while Table˜4 reports the errors achieved by the best model for each architecture. In addition, Figures˜10 and 11 display the predicted solutions and the corresponding error plots for the different PINO architectures. We can see that FC–PINO converges to the correct solution with minimal PDE residual.
@@ Conclusion
One long-term goal of the machine learning community is to enable progress on the millennium question concerning the finite-time blowup of the Navier–Stokes equation, where  [30; 31], already demonstrated that machine learning can identify a very important candidate for a self-similar blowup scenario for the axis-symmetric Euler equation with boundary [63; 64]. Generally, to upgrade numerical candidates to rigorous proofs, one uses rigorous analysis to show stability around the approximate profile with high precision. Further research in this area will have important implications for fluid blowup research.
@@ B.1.2 FC–Gram
We now describe the (Accelerated) FC–Gram method from [26] which is inspired from earlier FC–Gram methods [59; 23]. See [26] for more details and illustrations of FC–Gram.
@@
Each Gram polynomial is smoothly extended beyond the matching region using a blend-to-zero function defined over an extended interval consisting of: (i) a matching region of $d$ points, (ii) a continuation region of $c$ points, (iii) a zero region where the function vanishes. The extension is represented as a polynomial whose coefficients are precomputed by fitting the Gram polynomial in the matching region while ensuring decay in the zero region. This fitting uses oversampled grids and high-precision arithmetic to guarantee smoothness and numerical stability. The resulting functions are evaluated at the continuation nodes to form the columns of the continuation matrices $A_{\ell}\in\mathbb{R}^{c\times d}$ and $A_{r}\in\mathbb{R}^{c\times d}$.
@@ B.2 Tuning the Fourier Continuation Parameters
Tuning FC–Gram can be more challenging due to its localized and more complex construction. Parameters to tune include (1) the polynomial degree $d$, which controls the size of the local orthonormal basis and the number of matching points on each boundary segment, (2) the number of continuation points $c$, (3) the number of zero-blending points, (4) the number of extra points to improve numerical stability, (5) the oversampling factor to ensure accurate fine-grid construction, (6) the number of modes to reduce in the SVD computation to suppress the smallest singular values without significantly impacting resolution, and (7) the symbolic computation precision. However, after extensive tuning conducted in previous FC–Gram papers [59; 23; 26], fixed values of all these parameters except $d$ have been determined to provide an appropriate trade-off between overall computational cost and stability which works well in most settings. In particular, fixing $c=50$ works well across a range of $d$ values and generally does not require further tuning. Typically, fewer than $c=50$ extension points are needed to fully blend the signal to zero or the target constant, ensuring a gradual taper that avoids sharp discontinuities or artifacts.
@@
Overall, FC–Gram offers greater flexibility through its multiple tuning parameters but may require careful monitoring of local QR conditioning and noise overfitting. Its localized approach and blend-to-zero taper yield robust and clean derivative fields without manual filtering. For the 1D Burgers’ equation, Table˜8 shows that FC–Gram leads to high-precision solutions consistently across values of $d$, with minimal variations. This indicates that FC–Gram is very robust in this setting. With systematic validation and parameter sweeps, FC–Gram can be effectively tuned to deliver high-accuracy spectral continuation and differentiation. We emphasize that even without careful tuning of the parameter $d$ and keeping all other parameters fixed, we already obtained solutions with high accuracy.
@@
B.3 Properties and Comparison of FC–Legendre and FC–Gram
FC–Gram and FC–Legendre employ fundamentally different continuation strategies, each with unique pros and cons. As illustrated in Figure˜1, FC–Legendre constructs the extension by fitting a single Legendre polynomial of degree (2d1)(2d-1) spanning the entire extended interval between the left and right boundaries. In contrast, FC–Gram takes a more local approach, projecting boundary data onto lower-order polynomial bases of degree $d$ separately on each boundary and then smoothly blending the extensions to zero outside the domain. While both approaches yield good results with minimal tuning alongside FC–PINO, one can offer advantages over the other in high-precision settings.
Oscillation Attenuation and Stability.
FC–Gram achieves robustness through its blend-to-zero continuation strategy, which acts like a low-pass filter by smoothly damping boundary mismatches. This robustness is further strengthened by localized orthonormal Gram bases $Q_{\ell}$ and $Q_{r}$, constructed separately on the left and right boundary strips. These localized bases confine errors and high-frequency modes and noise within the extension region, preventing oscillations from contaminating the original domain signal. Consequently, spectral differentiation on the original domain interacts only with an attenuated boundary noise residual. This reduces the amplification of high-frequency components during differentiation and enhances spectral accuracy near boundaries. The localization also leads to better-conditioned QR factorizations of smaller Vandermonde matrices, reducing round-off errors and noise amplification during continuation.
@@ B.3 Properties and Comparison of FC–Legendre and FC–Gram
Simplicity.
FC–Legendre benefits from a simpler implementation, relying on straightforward global Legendre polynomial interpolation with few parameters to tune. FC–Gram, by contrast, requires a more complex setup involving Gram–Schmidt orthogonalization, oversampled least-squares fitting, and blend-to-zero tapering. These additional components provide greater robustness but introduce more parameters such as blend width, oversampling factor, and zero region size, that require careful tuning to optimize derivative accuracy and error confinement. However, after extensive tuning, we have found good parameters for most settings so that FC–Gram can be used with minimal adjustment (see Section˜B.2.2).
Computational Cost of the Extension Matrices.
FC–Gram generally requires higher-precision arithmetic than FC–Legendre due to fundamental differences in the polynomial basis construction. It builds its orthonormal basis numerically via a QR decomposition of Vandermonde-like matrices followed by least-squares projections. While this offers flexibility and localization, it is sensitive to rounding errors: poor scaling, near-dependent basis functions, and small matching intervals can degrade orthogonality, reducing spectral differentiation accuracy and increasing floating-point noise. Additionally, the matrix construction in FC–Gram is more computationally intensive than the analytic evaluation of Legendre polynomials in FC–Legendre. The latter benefits from analytically defined orthonormal polynomials computed with stable, optimized routines, resulting in better conditioning and less sensitivity to rounding errors. Both methods face conditioning challenges as $d$ grows, but FC–Gram is especially prone to amplified round-off errors. Consequently, although FC–Gram often offers greater noise robustness, it can require precision beyond float128 to ensure accurate, stable spectral continuation. We emphasize however, that these computational costs and precision requirements only apply for the construction of the extension matrices. For both FC–Legendre and FC–Gram, these matrices can be precomputed offline, and then reused in FC–PINO without the need for recomputation at a similar computation. Once precomputed, the costs of applying FC–Legendre and FC–Gram are very similar, and higher numerical precision is not necessary, unless solutions are desired with much higher precision.
Conditioning.
@@
Convergence of the derivatives
@@
For FC Gram, the continuation error on the physical interval consists of two parts, as demonstrated in [60; 61]. The dominant term is the interpolation error from matching $d$ boundary values, which converges at order $\mathcal{O}(N^{-d})$. In addition, there is a blending term caused by forcing the continuation to vanish smoothly outside the interval. This blending error does not formally converge with $N$, but can be made numerically negligible. When computing a derivative of order $k$, the interpolation error is reduced to order $\mathcal{O}(N^{-(d-k)})$, while the blending error remains bounded and does not affect the overall convergence rate.
@@
Now, FC–PINO is based on the Fourier neural operator, which enjoys universal approximation for sufficiently smooth operator mappings [5], i.e., one can make the approximation error $\|u_{\theta}-\mathcal{G}(f)\|$ arbitrarily small. By the Sobolev embedding theorem, with sufficient regularity, the error in the derivatives can also be controlled and kept arbitrarily small.
@@
In contrast, FC–Gram excels for signals that are less smooth, noisy, or contain discontinuities, as it provides greater robustness and effective localized error control. By constructing localized orthonormal polynomial bases on each boundary and smoothly blending the extension to zero, it confines boundary errors and high-frequency noise within the extension region, preventing their propagation into the main domain. This approach suppresses spurious oscillations near boundaries, resulting in cleaner derivatives without the need for manual filtering. Although FC–Gram is more computationally intensive and requires higher precision arithmetic, its QR factorizations performed on smaller local matrices produce better-conditioned continuation operators, which improve numerical stability and reduce sensitivity to round-off errors. Overall, it sacrifices some asymptotic accuracy to gain robustness and localized error confinement, making it especially well-suited for noisy or non-smooth problems.
@@ Appendix C Hyperparameters
We use a 2D FNO with 4 layers, each with 40 hidden channels and $(100,100)$ Fourier modes. We train for 60,000 epochs minimizing the MSE using Adam and the ReduceLROnPlateau scheduler with factor 0.5. We then conducted a grid search over the values of $\{64,100,128\}$ for the Fourier modes, $\{0.001,0.005,0.0001\}$ for the learning rate, and $\{500,1000,2000\}$ for the scheduler patience. For FC–Gram, we searched over $d$ in $\{3,4,5,6,7,8\}$, and for FC–Legendre, we searched over $c$ in $\{30,40,50,60,70,80\}$ and $d$ in $\{3,4,5,6,7,8\}$. Here, we employ ReLoBRaLo [37] to adaptively scale the loss coefficients. The configurations reported in Table˜10 correspond to the best model obtained for each approach.
@@ C.3 Navier–Stokes Equations
We use a 3D FNO with 4 layers, each with 60 hidden channels. We train for 60,000 epochs with a learning rate of 0.001 minimizing the MSE using Adam and the ReduceLROnPlateau scheduler with factor 0.5. We then conducted a grid search over learning rate in $\{0.001,0.005,0.0001\}$, scheduler patience in $\{500,1000,2000\}$, and Fourier modes in $\{24,32\}$. For FC–Gram, we searched over $d$ in $\{3,4,5,6\}$, and for FC–Legendre, we searched over $c$ in $\{20,30,40,50,60\}$ and $d$ in $\{3,4,5,6\}$. Here, we employ ReLoBRaLo [37] to adaptively scale the loss coefficients. The best configurations are reported in Table˜11.

### v3

@@ 4.2.2 Results
We compare the performance of FC–PINO to baseline models for solving a single instance of the 2D Burgers’ equation, where derivatives in the physics loss are computed using spectral differentiation. For each architecture, we performed a hyperparameter grid search, varying the number of Fourier modes, learning rate, scheduler patience, and the $(c,d)$ continuation parameters. The optimal hyperparameters for each architecture are provided in Section˜C.2, while Table˜3 reports the errors achieved by the best model for each architecture. FC–PINO performs very well, surpassing Pad–PINO by two orders of magnitude, while PINO Out–Pad and Standard PINO again fail to satisfy the governing equations.
@@ 4.3.2 Results
We compare the performance of FC–PINO to the baseline models for solving a single instance of the the Navier–Stokes equations, where derivatives in the physics loss are computed using spectral differentiation. For each architecture, we performed a hyperparameter grid search, varying the number of Fourier modes, learning rate, scheduler patience, and the $(c,d)$ continuation parameters.
The optimal hyperparameters for each architecture are provided in Section˜C.3, while Table˜4 reports the errors achieved by the best model for each architecture. In addition, Figures˜10 and 11 display the predicted solutions and the corresponding error plots for the different PINO architectures. We can see that FC–PINO converges to the correct solution with minimal PDE residual.
@@ Conclusion
One long-term goal of the machine learning community is to enable progress on the millennium question concerning the finite-time blowup of the Navier–Stokes equation, where (Wang et al., 2022, 2025b), already demonstrated that machine learning can identify a very important candidate for a self-similar blowup scenario for the axis-symmetric Euler equation with boundary (Luo and Hou, 2014a, b). Generally, to upgrade numerical candidates to rigorous proofs, one uses rigorous analysis to show stability around the approximate profile with high precision. Further research in this area will have important implications for fluid blowup research.
@@ A.1 Alternative FC-based PINO Architectures
Note that PINO In–FC and PINO Out–FC are mathematically equivalent, in the sense that they represent the same mapping and yield the same forward pass and derivatives in exact arithmetic. However, they induce different computational graphs for backpropagation, and therefore do not produce identical gradients in finite-precision arithmetic. In particular, even small discrepancies can be magnified when derivatives are computed via spectral differentiation and subsequently aggregated into the PDE residual loss. Consequently, although the two formulations are mathematically equivalent, the resulting backpropagated gradient can differ enough to steer the optimizer along substantially different trajectories, leading to different training dynamics and final model weights.
@@ B.1.2 FC–Gram
We now describe the (Accelerated) FC–Gram method from (Amlani and Bruno, 2016) which is inspired from earlier FC–Gram methods (Bruno et al., 2007; Albin and Bruno, 2011). See (Amlani and Bruno, 2016) for more details and illustrations of FC–Gram.
@@
Each Gram polynomial is smoothly extended beyond the matching region using a blending function defined over an extended interval consisting of: (i) a matching region of $d$ points, (ii) a continuation region of $c$ points, (iii) a region where the function vanishes. The extension is represented as a polynomial whose coefficients are precomputed by fitting the Gram polynomial in the matching region while ensuring decay in the zero region. This fitting uses oversampled grids and high-precision arithmetic to guarantee smoothness and numerical stability. The resulting functions are evaluated at the continuation nodes to form the columns of the continuation matrices $A_{\ell}\in\mathbb{R}^{c\times d}$ and $A_{r}\in\mathbb{R}^{c\times d}$.
@@ B.2 Tuning the Fourier Continuation Parameters
Tuning FC–Gram can be more challenging due to its localized and more complex construction. Parameters to tune include (1) the polynomial degree $d$, which controls the size of the local orthonormal basis and the number of matching points on each boundary segment, (2) the number of continuation points $c$, (3) the number of blending points, (4) the number of extra points to improve numerical stability, (5) the oversampling factor to ensure accurate fine-grid construction, (6) the number of modes to reduce in the SVD computation to suppress the smallest singular values without significantly impacting resolution, and (7) the symbolic computation precision. However, after extensive tuning conducted in previous papers (Bruno et al., 2007; Albin and Bruno, 2011; Amlani and Bruno, 2016), fixed values of all these parameters except $d$ have been determined to provide an appropriate trade-off between overall computational cost and stability which works well in most settings. In particular, fixing $c=50$ works well across a range of $d$ values and generally does not require further tuning. Typically, fewer than $c=50$ extension points are needed to fully blend the signal to the target constant, ensuring a gradual taper that avoids sharp discontinuities or artifacts.
@@
Overall, FC–Gram offers greater flexibility through its multiple tuning parameters but may require careful monitoring of local QR conditioning and noise overfitting. Its localized approach and blending taper yield robust and clean derivative fields without manual filtering. For the 1D Burgers’ equation, Table˜8 shows that FC–Gram leads to high-precision solutions consistently across values of $d$, with minimal variations. This indicates that FC–Gram is very robust in this setting. With systematic validation and parameter sweeps, FC–Gram can be effectively tuned to deliver high-accuracy spectral continuation and differentiation. We emphasize that even without careful tuning of the parameter $d$ and keeping all other parameters fixed, we already obtained solutions with high accuracy.
@@
B.3 Remark about Varying Resolution
The extensions obtained by Fourier continuation can depend heavily on the continuation parameters $c$ and $d$. For a given $(c,d)$ pair, the FC layer always uses $d$ boundary points on each side and produces an extension of length $c$, independently of the grid size $N$. As the resolution $N$ increases and the grid spacing decreases, the physical width of the boundary strips and extension region shrinks, so the continuation is built from a progressively smaller portion of the physical domain. Beyond a certain resolution, the strip of width $d$ can become too narrow in physical space to capture boundary layers or high frequency features, and the FC continuation may no longer reflect the true boundary behavior, even if the interior approximation remains accurate.
This behavior is analogous to the number of Fourier modes. If the number of modes is fixed while $N$ grows, the representation becomes increasingly band limited and eventually fails to resolve the relevant dynamics. In the same way, fixed values of the continuation parameters $c$ and $d$ can become insufficient at high resolution, which can also prove problematic when the model is trained with losses at one resolution and evaluated at a much finer resolution in super-resolution or cross-resolution settings. In practice, one can choose $(c,d)$ that are robust over a target range of resolutions and further stabilize performance with multi-resolution training or progressive schedules that increase the number of Fourier modes together with the continuation parameters. The incremental FNO strategy (George et al., 2022), where the resolution and number of modes is increased progressively, can also be extended by progressively increasing the values of the Fourier continuation parameters $c$ and $d$.
B.4 Properties and Comparison of FC–Legendre and FC–Gram
FC–Gram and FC–Legendre employ fundamentally different continuation strategies, each with unique pros and cons. As illustrated in Figure˜1, FC–Legendre constructs the extension by fitting a single Legendre polynomial of degree (2d1)(2d-1) spanning the entire extended interval between the left and right boundaries. In contrast, FC–Gram takes a more local approach, projecting boundary data onto lower-order polynomial bases of degree $d$ separately on each boundary and then smoothly blending the extensions outside the domain. While both approaches yield good results with minimal tuning alongside FC–PINO, one can offer advantages over the other in high-precision settings.
B.4.1 Oscillation Attenuation and Stability
FC–Gram achieves robustness through its blending continuation strategy, which acts like a low-pass filter by smoothly damping boundary mismatches. This robustness is further strengthened by localized orthonormal Gram bases $Q_{\ell}$ and $Q_{r}$, constructed separately on the left and right boundary strips. These localized bases confine errors and high-frequency modes and noise within the extension region, preventing oscillations from contaminating the original domain signal. Consequently, spectral differentiation on the original domain interacts only with an attenuated boundary noise residual. This reduces the amplification of high-frequency components during differentiation and enhances spectral accuracy near boundaries. The localization also leads to better-conditioned QR factorizations of smaller Vandermonde matrices, reducing round-off errors and noise amplification during continuation.
@@ B.4.1 Oscillation Attenuation and Stability
B.4.2 Simplicity
FC–Legendre benefits from a simpler implementation, relying on straightforward global Legendre polynomial interpolation with few parameters to tune. FC–Gram, by contrast, requires a more complex setup involving Gram–Schmidt orthogonalization, oversampled least-squares fitting, and blending tapering. These additional components provide greater robustness but introduce more parameters such as blend width, oversampling factor, and zero region size, that require careful tuning to optimize derivative accuracy and error confinement. However, after extensive tuning, we have found good parameters for most settings so that FC–Gram can be used with minimal adjustment (see Section˜B.2.2).
B.4.3 Computational Cost of the Extension Matrices
FC–Gram generally requires higher-precision arithmetic than FC–Legendre due to fundamental differences in the polynomial basis construction. It builds its orthonormal basis numerically via a QR decomposition of Vandermonde-like matrices followed by least-squares projections. While this offers flexibility and localization, it is sensitive to rounding errors: poor scaling, near-dependent basis functions, and small matching intervals can degrade orthogonality, reducing spectral differentiation accuracy and increasing floating-point noise. Additionally, the matrix construction in FC–Gram is more computationally intensive than the analytic evaluation of Legendre polynomials in FC–Legendre. The latter benefits from analytically defined orthonormal polynomials computed with stable, optimized routines, resulting in better conditioning and less sensitivity to rounding errors. Both methods face conditioning challenges as $d$ grows, but FC–Gram is especially prone to amplified round-off errors. Consequently, although FC–Gram often offers greater noise robustness, it can require precision beyond float128 to ensure accurate, stable spectral continuation.
We emphasize however, that these computational costs and precision requirements only apply for the construction of the extension matrices. For both FC–Legendre and FC–Gram, these matrices can be precomputed offline, and then reused in FC–PINO without the need for recomputation at a similar computation. Once precomputed, the costs of applying FC–Legendre and FC–Gram are very similar, and higher numerical precision is not necessary, unless solutions are desired with much higher precision.
B.4.4 Conditioning
@@
B.4.5 Convergence of the derivatives
Let $u_{\theta}$ denote the output of the PINO model, $\widetilde{u}_{\theta}$ its FC extension, and $u$ the true solution.
In FC–PINO, spectral differentiation is performed on the continuation $\widetilde{u}_{\theta}$, so the error made on the derivatives can be decomposed as
@@ B.4.5 Convergence of the derivatives
For FC Gram, the continuation error on the physical interval consists of two parts (Lyon, 2009; Lyon and Bruno, 2010). The dominant term is the interpolation error from matching $d$ boundary values, which converges at order $\mathcal{O}(N^{-d})$. In addition, there is a blending term caused by forcing the continuation to vanish smoothly outside the interval. This blending error does not formally converge with $N$, but can be made numerically negligible. When computing a derivative of order $k$, the interpolation error is reduced to order $\mathcal{O}(N^{-(d-k)})$, while the blending error remains bounded and does not affect the overall convergence rate.
@@
Now, FC–PINO is based on the Fourier neural operator, which enjoys universal approximation for sufficiently smooth operator mappings (Kovachki et al., 2021a), i.e., one can make the approximation error $\|u_{\theta}-\mathcal{G}(f)\|$ arbitrarily small. By the Sobolev embedding theorem, with sufficient regularity, the error in the derivatives can also be controlled and kept arbitrarily small.
@@ B.4.6 Summary
In contrast, FC–Gram excels for signals that are less smooth, noisy, or contain discontinuities, as it provides greater robustness and effective localized error control. By constructing localized orthonormal polynomial bases on each boundary and smoothly blending the extension, it confines boundary errors and high-frequency noise within the extension region, preventing their propagation into the main domain. This approach suppresses spurious oscillations near boundaries, resulting in cleaner derivatives without the need for manual filtering. Although FC–Gram is more computationally intensive and requires higher precision arithmetic, its QR factorizations performed on smaller local matrices produce better-conditioned continuation operators, which improve numerical stability and reduce sensitivity to round-off errors. Overall, it sacrifices some asymptotic accuracy to gain robustness and localized error confinement, making it especially well-suited for noisy or non-smooth problems.
@@ Appendix C Hyperparameters
We use a 2D FNO with 4 layers, each with 40 hidden channels and $(100,100)$ Fourier modes. We train for 60,000 epochs minimizing the MSE using Adam and the ReduceLROnPlateau scheduler with factor 0.5. We then conducted a grid search over the values of $\{64,100,128\}$ for the Fourier modes, $\{0.001,0.005,0.0001\}$ for the learning rate, and $\{500,1000,2000\}$ for the scheduler patience. For FC–Gram, we searched over $d$ in $\{3,4,5,6,7,8\}$, and for FC–Legendre, we searched over $c$ in $\{30,40,50,60,70,80\}$ and $d$ in $\{3,4,5,6,7,8\}$. Here, we employ ReLoBRaLo to adaptively scale the loss coefficients. The configurations reported in Table˜10 correspond to the best model obtained for each approach.
@@ C.3 Navier–Stokes Equations
We use a 3D FNO with 4 layers, each with 60 hidden channels. We train for 60,000 epochs with a learning rate of 0.001 minimizing the MSE using Adam and the ReduceLROnPlateau scheduler with factor 0.5. We then conducted a grid search over learning rate in $\{0.001,0.005,0.0001\}$, scheduler patience in $\{500,1000,2000\}$, and Fourier modes in $\{24,32\}$. For FC–Gram, we searched over $d$ in $\{3,4,5,6\}$, and for FC–Legendre, we searched over $c$ in $\{20,30,40,50,60\}$ and $d$ in $\{3,4,5,6\}$. Here, we employ ReLoBRaLo to adaptively scale the loss coefficients. The best configurations are reported in Table˜11.